# A dimension reduction approach to edge weight estimation for use in spatial models

### v1

@@ 4 Illustrations
Within this section we evaluate the performance of our approach under a variety of settings and in comparison to alternative methods. We consider both simulated and real-world data examples.
4.1 Simulations
We begin with three simulation studies. In the first we evaluate approximate confidence interval coverage rates when the model is correctly specified. In the second, we consider how to select the number of basis functions when implementing our method. The third simulation assesses model performance under misspecification, evaluating the Kullback-Leibler divergence between the true data generating distribution and the estimated distribution.
Sim 1: Approximate confidence interval coverage
We defined $10\times 10$ and $20\times 20$ lattices, and obtained the LGL eigenvectors for each. We selected $k$ eigenvectors and simulated $\eta\sim N_{k}(\textbf{0},\Phi)$, where $\Phi$ is a diagonal matrix with first entry (corresponding to the variance of the “intercept” eigenvector) equal to 0.5 and all other diagonal entries equal to 25. We then generated $n$ samples from $N_{p}(\textbf{0},\Sigma(\eta))$ where $\Sigma(\eta)$ is defined by Equations 4 and 6, using a Matérn covariance with $\nu=3/2$ and fixing $\sigma^{2}=1$. This was done for each of $k\in\{10,30\}$ and $n\in\{1,10,50\}$. We obtained the MLE $\hat{\eta}$ and calculated approximate 90% confidence intervals for $\eta$ using Fisher’s approximation. The simulation was repeated ten times, with coverage rates provided in Table 1.
$n=1$ $n=10$ $n=50$
$p=100,\;k=10$ 0.880 ($\pm$ 0.039) 0.910 ($\pm$ 0.023) 0.870 ($\pm$ 0.040)
$p=100,\;k=30$ 0.804 ($\pm$ 0.040) 0.867 ($\pm$ 0.027) 0.927 ($\pm$ 0.015)
$p=400,\;k=10$ 0.920 ($\pm$ 0.025) 0.910 ($\pm$ 0.023) 0.910 ($\pm$ 0.031)
$p=400,\;k=30$ 0.907 ($\pm$ 0.018) 0.900 ($\pm$ 0.021) 0.883 ($\pm$ 0.019)
Table 1: Coverage rates of approximate 90% confidence intervals. Monte Carlo standard error in parentheses.
In general coverage rates are appropriate, the exception being under the settings $k=30$ and $n=1$ over a $10\times 10$ lattice. On top of having lower than expected coverage, Newton’s method diverged in one of the ten simulations under these settings, indicating that one observation of a spatial process over a graph with $p=100$ nodes may not be enough to inform the estimation of $\Sigma$ using $k=30$ basis functions. In general however, we feel that inference based on Fisher’s approximation to the distribution of the MLE is appropriate. We investigate choice of $k$ in the following simulation.
Sim 2: Number of basis functions
Once again, we defined $10\times 10$ and $20\times 20$ lattices obtaining LGL eigenvectors for each. Instead of simulating edge weights directly from Equation 6 we generate an edge weights matrix W from an ICAR model using the line graph adjacency matrix to define the associations between edges. Figure 8 depicts sampled edge weights over a $20\times 20$ lattice. The spatial dependence in the simulated edge weights themselves means that W can be well approximated using our method. Larger numbers of basis functions will improve this approximation, but at the cost of model complexity and potential computational instability. The simulated W was used to define a covariance matrix according to Equation 4 with $\nu=3/2$ and $\sigma^{2}=1$. Independent samples were generated from $N_{p}(\textbf{0},\Sigma)$ and stored in $n\times p$ matrix Y for $n\in\{1,10,50\}$. For each simulated Y we fit a GDEF model using Matérn covariance with $\nu=3/2$ using $k\in\{10,20,...,100\}$ basis functions and evaluated AIC and BIC for each of the 10 model fits, with lower values of
AIC $=-2\ell(\eta;\textbf{Y})+2k$ (12)
BIC $=-2\ell(\eta;\textbf{Y})+k\text{ log}(np)$
begin selected. The simulation was repeated 10 times. Table 2 contains information regarding the AIC- and BIC-preferred choices of $k$.
Figure 8: Simulated W on left, with $\hat{\textbf{W}}$ on right (estimated from $n=1$ observations and using $k=20$ basis functions). Approximation “smooths” the true edge weights.
@@ 4.1 Simulations
Across simulation settings, we see that models with higher $k$ perform better on the larger grid and with bigger sample sizes. This is unsurprising as larger amounts of data should be expected to support more complicated models. AIC and BIC are considered not because they are objectively the best tools for model selection, but because they provide a likelihood-based starting point for the question of how to choose the number of eigenvectors when implementing our method. BIC is more conservative than AIC due to a larger model size penalty. The two criteria also differ in that AIC is calibrated for improving predictive accuracy, while BIC may be viewed as favoring the model that best approximates the marginal distribution of the data (Gelman et al., 2014). The fact that the two criteria frequently prefer different values of $k$ provides some intuition for the range on model size that may be appropriate. Using the information in Table 2, a reasonable starting point is to select $k=\sqrt{np}$ when the model reasonably approximates the true data generating process, as was the case within this simulation. Optimal choice of $k$ will also be contingent on the nature of the spatial process itself, as evidence of more complicated patterns of spatial dependence may indicate that larger numbers of basis functions should be used. Table 3 contains the rates at which the Newton-Raphson algorithm used to estimate $\hat{\eta}$ failed to converge across different simulation settings, with the majority of instances occurring for large $k$ and $n=1$. When computational instability is encountered, refitting with smaller $k$ will frequently resolve the issue; the possibility of instability due to significant model misspecification should however be considered.
$k=40$ $k=50$ $k=60$ $k=70$ $k=80$ $k=90$ $k=100$
$p=100,\;n=1$ 0.3 0.7 0.8 1.0 1.0 1.0 1.0
$p=100,\;n=10$ 0.0 0.0 0.0 0.0 0.0 0.2 0.1
$p=100,\;n=50$ 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Table 3: Frequency with which Newton-Raphson algorithm failed to converge. Estimation was stable for all $k<40$ and for all settings over the $20\times 20$ lattice.
Sim 3: Model performance under misspecification
The previous simulation evaluated the performance of the GDEF model for different values of $k$ when the true data generating process was a GDEF model with an edge weights matrix from outside of the parameter space characterized by Equation 6. We now consider the performance of the GDEF model when data comes from a distinct nonstationary covariance model. Given a nonstationary spatial process $Z(\textbf{s})$ observed at $\textbf{s}\in\mathbb{R}^{2}$ the spatial deformation approach of Sampson and Guttorp (1992) defines the process’ covariance such that
$\text{Cov}(Z(\textbf{s},\textbf{s}^{\prime}))=\sigma^{2}\rho(||g(\textbf{s})-g(\textbf{s}^{\prime})||)$ (13)
where $g(\cdot)$ is a transformation from the geographic location s to its location $g(\textbf{s})$ in a “deformed” space wherein covariance is stationary and isotropic.
@@
Inspired by this approach, we simulate data by first defining a $15\times 15$ lattice. We then deform the grid by randomly selecting ten “deformation points” $d_{1:10}\in(1,15)^{2}$ before sequentially shrinking or expanding the grid towards or away from these locations, as illustrated by Figure 9. Let $\textbf{D}^{D}$ be the $225\times 225$ matrix of Euclidean distances between the nodes of the deformed lattice. Applying a Matérn correlation function to $\textbf{D}^{D}$ results in a covariance matrix that is nonstationary with respect to D, the matrix of Euclidean distances between the original grid point locations. We generate $n$ observations from the following model
$\textbf{y}_{i}$ $=\beta_{0}\textbf{1}+\textbf{z}_{i}+\epsilon_{i}$ (14)
$\textbf{z}_{i}$ $\sim N_{p}(\textbf{0},\sigma^{2}\Phi)$
$\epsilon_{i}$ $\sim N_{p}(\textbf{0},\tau^{2}\textbf{I}_{p})$
where $\textbf{z}_{i}$ is a vector of spatial random effects with covariance matrix $\Phi$ defined using a Matérn correlation function with smoothness $\nu=5/2$
$\Phi=\left(1+\frac{\sqrt{5}}{\rho}\textbf{D}^{D}+\frac{5}{3\rho^{2}}\{\textbf{D}^{D}\}^{\circ 2}\right)\odot\text{exp}\left(\frac{\sqrt{5}}{\rho}\textbf{D}^{D}\right)$ (15)
@@
We set $\beta_{0}=0$, $\sigma^{2}=0.9$, $\tau^{2}=0.1$, $\rho=3$. We fit the GDEF and other models to the simulated Y as illustrated by Figure 10, which depicts a deformed lattice, samples of $\textbf{y}_{i}$ produced by that deformation, and an estimated edge weights matrix.
Figure 9: Deformed lattice obtained by repeatedly expanding and contracting grid around random locations. Each simulation replicate used a uniquely generated deformation.
@@
We randomly generated 125 unique grid deformations. For each deformation we sampled Y as described above for $n\in\{1,2,3,5,10,25,50\}$ and fit Equation 14 while defining the covariance of the random effects $\textbf{z}_{i}$ using the following six submodels: a GDEF model with $\nu=3/2$ and $k=15$ basis functions, a GDEF model using the double exponential covariance function $(\nu=\infty)$ and $k=15$ basis functions, the spatial deformation approach of Sampson and Guttorp (1992) (S&G) as implemented by the deform R package (Youngman, 2023), a simple Matérn covariance model with $\nu=5/2$ and using the Euclidean distances between non-deformed lattice nodes, the method presented in Hughes and Haran (2013) (H&H) for modeling random effects in areal data using 100 basis functions, and a standard CAR model using an unweighted adjacency matrix. The spatial deformation approach and the stationary Matérn covariance model treat the data as though it were point-indexed, while all other approaches utilize only the lattice’s graphical structure; note however, that the true data generating process exists in continuous space.
@@
$\text{KL}(\Sigma,\hat{\Sigma})=\frac{1}{2}\left(\text{log}\frac{|\hat{\Sigma}|}{|\Sigma|}+\text{tr}\left(\hat{\Sigma}^{-1}\Sigma)\right)-p\right)$ (16)
@@
Table 4 depicts the average KL divergence between the true distributions and the distributions estimated by the six models. Note that the true distribution is different for each of the 125 replicates which used a unique deformation. We also consider the frequency with which each model had the smallest KL divergence under the same settings. As can be seen, the GDEF model (for both choices of $\nu$) performs well across sample sizes, with the simple Matérn model for $n=1$ being the only alternative that was frequently competitive in this simulation. While potentially surprising that the simple Matérn covariance model using the non-deformed distances outclasses many of the more complicated alternatives under some settings, it is clear that its relative performance is strongest for small $n$; the true data generating process also used a Matérn covariance function with $\nu=5/2$, and a visual assessment of simulation output suggests that the simple Matérn performs best when the randomly generated deformation exhibits less dramatic patterns of warping.
GDEF 3/2 GDEF $\infty$ S&G Matérn 5/2 H&H Std. CAR
$n=1$ 42.54 (0.48) 46.66 (0.192) - 45.22 (0.328) 96.06 76.74
$n=2$ 28.2 (0.576) 29.61 (0.376) 73.6 42.35 (0.048) 144.81 79.92
$n=3$ 24.24 (0.6) 24.74 (0.384) 40.29 40.87 (0.016) 77.6 81.14
$n=5$ 21.93 (0.456) 21.94 (0.536) 32.33 (0.008) 39.75 55.57 83.24
$n=10$ 20.6 (0.304) 20.4 (0.688) 29.54 (0.008) 39.13 48.4 83.95
$n=25$ 19.83 (0.248) 19.56 (0.728) 28.53 (0.024) 38.65 46.31 84.45
$n=50$ 19.55 (0.24) 19.25 (0.744) 28.3 (0.016) 38.36 45.9 85.08
@@
Table 4: Average $\text{KL}(\Sigma,\hat{\Sigma})$ for each model and number of samples. In parentheses is the frequency across simulations that a given model produced the smallest KL divergence for a sample of size $n$. The GDEF models perform best across all sample sizes.
@@
Note that the spatial deformation approach (S&G) cannot be fit for $n=1$. This highlights one of the key advantages of our approach. Much of the literature regarding the spatial deformation framework requires that the spatial process be observed multiple times, and in some implementations requires that $n>p$. However, it is extremely common for spatial analyses to be conducted on datasets with little or no replication. Whether the deformation function is estimated using thin plate splines as in Sampson and Guttorp (1992) or a Gaussian process prior as in Schmidt and O’Hagan (2003), the inherent flexibility of the framework requires considerable amounts of data (or strong prior assumptions) to avoid over fitting. As seen in this and previous simulations, the GDEF model using the basis function representation of W as presented in this paper can effectively estimate complicated patterns of spatial dependence for areal data and performs well in a setting where the true spatial process exists in continuous rather than discrete space.
@@ 4.2 Example: Mercer and Hall wheat yield data
We conclude this section with an analysis of the classic wheat yield dataset from Mercer and Hall (1911) and which is available in the spData R package. The version of the data used in the package was taken from Cressie (1993). Mercer and Hall considered a one acre parcel of farmland that was divided into 500 rectangular plots and planted with wheat. Each plot received the same treatment and soil quality was perceived to be essentially uniform at the time of planting. At harvest, the yield of each plot in pounds of grain was recorded. Figure 11 depicts the yields for each plot, which were arranged in 20 rows each containing 25 plots. Mercer and Hall discuss the variability present in the data and explore the possibility of location-based effects on yield. Noting evidence of spatial correlation, the wheat yield data was analyzed in Whittle (1954), which first introduced the simultaneous autoregressive (SAR) model, and in Besag (1974), which first introduced the CAR model. The authors of both papers found the fit of their models to the wheat yield data unsatisfactory. Looking at Figure 11, one notes that certain columns (4,7 and 10) have larger average yields than others and that within-column correlation appears to be stronger than within-row correlation. By modeling the data using fixed row and column effects Cressie (1993) found that a stationary and isotropic kriging model with plot locations given by centroids resulted in satisfactory fit, though it is worth considering whether the inclusion of row and column effects in a model for this data provide interpretive value or merely serve to improve the fit of an otherwise poorly specified model.
Figure 11: The Mercer Hall wheat yield data
We propose modeling this data according to Equation 14 with the covariance of the random effects defined by a GDEF model with smoothness $\nu=3/2$ and $k=20$ basis functions. To allow for a distinction between within-column and within-row dependency, we split the intercept vector $\textbf{v}_{1}$ of our basis into two vectors, $\textbf{v}_{1}^{row}$ and $\textbf{v}_{1}^{col}$, where $v_{i1}^{r}=1$ if edge $i$ connects two nodes in the same row and is 0 otherwise, with $\textbf{v}_{1}^{col}$ defined similarly for columns. We obtain maximum likelihood estimates $\hat{\eta}$, $\hat{\sigma^{2}}$, $\hat{\tau^{2}}$ and $\hat{\beta_{0}}$ which are reported in Table 5. Negative $\hat{\eta_{1}}^{row}$ and positive $\hat{\eta_{1}}^{col}$ indicates much stronger within-column than within-row correlations. Rather than reporting the other coefficients in $\hat{\eta}$ which are individually less interpretable, Figure 12 depicts log($\hat{\textbf{w}}$) over the entire network; there are three regions within the spatial domain with larger edges indicating stronger spatial correlation between nodes in those locations.
Figure 12: Estimated log edge weights
@@
We define $\hat{\textbf{z}}=\left(\hat{\Phi}^{-1}/\hat{\sigma^{2}}+\textbf{I}_{p}/\hat{\tau^{2}}\right)^{-1}(\textbf{y}-\hat{\beta_{0}}\textbf{1})/\hat{\tau^{2}}$ and provide a plot of $\hat{\textbf{z}}$ in Figure 13 alongside $\textbf{y}-\hat{\beta_{0}}\textbf{1}$. Moran’s I (Moran, 1950) is a test statistic indicating the presence or absence of spatial autocorrelation. Moran’s I for the residuals of our model $\epsilon=\textbf{y}-\hat{\beta_{0}}\textbf{1}-\hat{\textbf{z}}$ is 0.022, with a p-value of 0.469, indicating a lack of spatial correlation in $\epsilon$. The Shapiro-Wilk normality test (Shapiro and Wilk, 1965) on the residuals results in a p-value of 0.262 indicating that the normality assumption for $\epsilon$ is not violated. We find that the GDEF model fits the wheat yield data well, effectively representing the distinctive and irregular patterns of spatial dependence (seen especially in the columns of the data) in an uncontrived manner.
Figure 13: Estimate of z (left) which acts as a smoothed version of the centered data (right)

### v2

@@ 4 Illustrations
In this section we evaluate the performance of the GDEF model using the LGL eigenvector basis in both simulated and real-world data settings. We begin with two simulation studies. In the first, we consider how to select the number of basis functions when implementing our method. The second simulation assesses model performance under misspecification, evaluating the Kullback-Leibler divergence between the true data generating distribution and the estimated distribution. The Supplemental material contains an additional simulation study demonstrating coverage rates of the approximate confidence intervals obtained using Equation 8, as well as additional details on the two simulations presented here.
4.1 Simulation 1: Number of basis functions
We defined $\times 10$ 10 and $\times 20$ 20 lattices obtaining LGL eigenvectors for each. Instead of simulating edge weights directly from Equation 6, we generate an edge weights matrix W from an ICAR model using the line graph adjacency matrix to define the associations between edges. Figure 3 of the supplemental material for this article depicts one such sampled set of edge weights over a $\times 20$ 20 lattice. The spatial dependence in the simulated edge weights themselves means that W can be well approximated using our method. Larger numbers of basis functions will improve this approximation, but at the cost of model complexity and potential computational instability. The simulated W was used to define a covariance matrix according to Equation 4 with $\nu=3/2$ and $\sigma^{2}=1$. Independent samples were generated from $N_{p}(\textbf{0},\Sigma)$ and stored in $n\times p$ matrix Y for $n\in\{1,10,50\}$. For each simulated Y we fit a GDEF model using Matérn covariance with $\nu=3/2$ using $k\in\{10,20,...,100\}$ basis functions and evaluated AIC and BIC for each of the 10 model fits. The simulation was repeated 10 times. Table 1 contains information regarding the AIC- and BIC-preferred choices of $k$.
@@ 4.1 Simulation 1: Number of basis functions
Across simulation settings, we see that models with higher $k$ perform better on the larger grid and with bigger sample sizes. Using the information in Table 1, a potential starting point is to select $k=\sqrt{np}$ when the model reasonably approximates the true data generating process, as was the case within this simulation. Optimal choice of $k$ will also be contingent on the nature of the spatial process itself, as evidence of more complicated patterns of spatial dependence may indicate that larger numbers of basis functions should be used.
4.2 Simulation 2: Model performance under misspecification
We compare the performance of the GDEF model with others in a setting where the GDEF model is misspecified. Data is generated by the nonstationary spatial process obtained by the spatial deformation approach of Sampson and Guttorp, (1992). This defines the process covariance
$\text{Cov}(z(\textbf{s}),z(\textbf{s}^{\prime}))=\sigma^{2}\rho(||g(\textbf{s})-g(\textbf{s}^{\prime})||),$ (9)
@@ 4.2 Simulation 2: Model performance under misspecification
where $g(\cdot)$ is a transformation from the geographic location s to its location $g(\textbf{s})$ in a “deformed” space wherein covariance is stationary and isotropic. Data are simulated by first defining a $\times 15$ 15 lattice. We then generate random deformations of the lattice by contracting or expanding the grid towards or away from randomly sampled locations. Let $\textbf{D}^{D}$ be the $\times 225$ 225 matrix of Euclidean distances between the nodes of the deformed lattice. Applying a Matérn correlation function to $\textbf{D}^{D}$ results in a covariance matrix that is nonstationary with respect to D, the matrix of Euclidean distances between the original grid point locations. We generate $n$ observations from the following model
$\textbf{y}_{t}$ $=\beta_{0}\textbf{1}+\textbf{z}_{t}+\epsilon_{t},$ (10)
$\textbf{z}_{t}$ $\sim N_{p}(\textbf{0},\sigma^{2}\rho_{5/2}),$
$\epsilon_{t}$ $\sim N_{p}(\textbf{0},\tau^{2}\textbf{I}_{p}),$
@@
where $\textbf{z}_{t}$ is a vector of spatial random effects with correlation matrix $\rho_{5/2}$ defined using a Matérn correlation function with smoothness $\nu=5/2$. We set $\beta_{0}=0$, $\sigma^{2}=0.9$, $\tau^{2}=0.1$ and a range parameter of 333. We fit the GDEF and other models to the simulated data Y as illustrated by Figure 6, which depicts a deformed lattice, samples of $\textbf{y}_{t}$ produced by that deformation, and an estimated edge weights matrix.
@@
We randomly generated 125 unique grid deformations. For each deformation we sampled Y as described above for $n\in\{1,2,3,5,10,25,50\}$ and fit Equation 10 while defining the covariance of the random effects $\textbf{z}_{t}$ using the following five submodels: a GDEF model with $\nu=3/2$ and $k=15$ basis functions, the spatial deformation approach of Sampson and Guttorp, (1992) (S&G) as implemented by the deform R package (Youngman,, 2023), a simple Matérn covariance model with $\nu=5/2$ and using the Euclidean distances between non-deformed lattice nodes, the method presented in Hughes and Haran, (2013) (H&H) for modeling random effects in areal data using 50 basis functions, and a standard CAR model using an unweighted adjacency matrix. The spatial deformation approach and the stationary Matérn covariance model treat the data as though it were point-indexed, while all other approaches utilize only the lattice’s graphical structure; note however, that the true data generating process exists in continuous space.
@@
$\text{KL}(\Sigma,\hat{\Sigma})=\frac{1}{2}\left(\text{log}\frac{|\hat{\Sigma}|}{|\Sigma|}+\text{tr}\left(\hat{\Sigma}^{-1}\Sigma)\right)-p\right).$ (11)
@@
Table 2 depicts the average KL divergence between the true distributions and the distributions estimated by the five models. Note that the true distribution is different for each of the 125 replicates which used a unique deformation. We also consider the frequency with which each model had the smallest KL divergence under the same settings. As can be seen, the GDEF model performs well across sample sizes, though the simple Matérn and CAR models frequently produce competitive estimates of $\Sigma$ when $n=1$. It is intuitive that simpler models would perform better in contexts with limited data available, and visual inspection of simulation output suggests that the simple Matérn and CAR models perform best in comparison to the GDEF approach when the randomly generated deformation exhibits less dramatic patterns of warping. The GDEF model is the best performing in settings with 10 or fewer replicates, though in settings with $n=25$ and $n=50$ it is surpassed on average by the S&G approach, which under the settings of this simulation study is closest to the true data generating model. The fact that the GDEF model consistently outperforms S&G for $n\leq 10$ suggests that its internal structure provides some degree of stability and regularization when estimating complex covariance structures using limited numbers of replicates.
GDEF S&G Matérn 5/2 H&H Std. CAR
$n=1$ 42.54 (0.552) - 45.22 (0.24) 70.85 43.4 (0.208)
$n=2$ 28.2 (0.928) 73.6 42.35 (0.04) 69.35 42.28 (0.032)
$n=3$ 24.24 (0.984) 40.29 40.87 (0.016) 75.78 41.88
$n=5$ 21.93 (0.992) 32.33 (0.008) 39.75 58.58 41.67
$n=10$ 20.6 (0.896) 26.54 (0.104) 39.13 51.75 41.56
$n=25$ 19.83 (0.376) 19.09 (0.624) 38.65 50.10 41.44
$n=50$ 19.55 (0.192) 17.02 (0.808) 38.36 49.7 41.39
@@
Table 2: Average $\text{KL}(\Sigma,\hat{\Sigma})$ for each model and number of samples. In parentheses is the frequency across simulations that a given model produced the smallest KL divergence for a sample of size $n$. The GDEF model performs well across settings, and especially with limited numbers of replicates.
@@
Note that the spatial deformation approach (S&G) cannot be fit for $n=1$. This highlights one of the key advantages of our approach. Much of the literature regarding the spatial deformation framework requires that the spatial process be observed multiple times, and in some implementations requires that $n>p$. However, it is extremely common for spatial analyses to be conducted on datasets with little or no replication. Whether the deformation function is estimated using thin plate splines as in Sampson and Guttorp, (1992) or a Gaussian process prior as in Schmidt and O’Hagan, (2003), the inherent flexibility of the framework requires considerable amounts of data (or strong prior assumptions) to avoid over-fitting. As seen here the GDEF model using the basis function representation of W as presented in this paper can effectively estimate complicated patterns of spatial dependence for areal data and performs well in a setting where the true spatial process exists in continuous rather than discrete space.
@@ 4.3 Example: Mercer and Hall wheat yield data
We conclude this section with an analysis of the classic wheat yield dataset from Mercer and Hall, (1911) and which is available in the spData R package. The version of the data used in the package was taken from Cressie, (1993). Mercer and Hall, (1911) considered a one acre parcel of farmland that was divided into 500 rectangular plots and planted with wheat. Each plot received the same treatment and soil quality was perceived to be essentially uniform at the time of planting. At harvest, the yield of each plot in pounds of grain was recorded. Figure 7 depicts the yields for each plot, which were arranged in 20 rows each containing 25 plots. Mercer and Hall discuss the variability present in the data and explore the possibility of location-based effects on yield. Noting evidence of spatial correlation, the wheat yield data was analyzed in Whittle, (1954), which first introduced the simultaneous autoregressive (SAR) model, and in Besag, (1974), which first introduced the CAR model. The authors of both papers found the fit of their models to the wheat yield data unsatisfactory. Columns (4,7 and 10) in Figure 7 have larger average yields than others and the within-column correlation appears to be stronger than within-row correlation. By modeling the data using fixed row and column effects (Cressie,, 1993) found that a stationary and isotropic kriging model with plot locations given by centroids resulted in satisfactory fit, though it is worth considering whether the inclusion of row and column effects in a model for this data provide interpretive value or merely serve to improve the fit of an otherwise poorly specified model.
Figure 7: The Mercer Hall wheat yield data. Each cell corresponds to an agricultural plot; response variable is pounds of grain at harvest.
[…]
Figure 8: Estimated log edge weights (left) for the Mercer-Hall data which in turn define the covariance structure of z, the “de-noised” version of the centered data (right).
@@
We define $\hat{\textbf{z}}=\left(\hat{\Phi}^{-1}/\hat{\sigma^{2}}+\textbf{I}_{p}/\hat{\tau^{2}}\right)^{-1}(\textbf{y}-\hat{\beta_{0}}\textbf{1})/\hat{\tau^{2}}$ and provide a plot of $\hat{\textbf{z}}$ in the right half of Figure 8. Note that regions where the estimate of z are smoother correspond to larger estimated edge weights. Moran’s I (Moran,, 1950) is a test statistic indicating the presence or absence of spatial autocorrelation. Moran’s I for the estimated residuals of our model $\hat{\epsilon}=\textbf{y}-\hat{\beta_{0}}\textbf{1}-\hat{\textbf{z}}$ is 0.022, with a p-value of 0.460, indicating a lack of spatial correlation in $\hat{\epsilon}$. The Shapiro-Wilk normality test (Shapiro and Wilk,, 1965) on the estimated residuals results in a p-value of 0.262 indicating that the normality assumption for $\epsilon$ is not violated.
To assess the out-of-sample predictive performance of the GDEF model using the LGL eigenvector basis on this data set in comparison to existing methods, we perform ten-fold spatial cross validation, partitioning the 500 plot locations into 20 five-by-five blocks of plots, then splitting the 20 blocks into ten sets of two, resulting in a total of 50 plots per split. In addition to the GDEF model described above, we fit each of the following models to the test-training splits: a Matérn covariance model with $\nu=5/2$ and distances computed using plot centroids, the spatial random effects model of Hughes and Haran, (2013) using 50 basis functions, a standard CAR model, and a CAR model with two correlation parameters, one for each of the within-row and within-column dependencies, a modeling consideration similar in nature to the split we made of the intercept column in the basis representation of the GDEF edge weights. For each model and test-training split, we compute the root mean square error (RMSE) and mean absolute error (MAE), and produce 95% prediction intervals for all elements of the test set. In addition to our comparison of predictive performance, we wish to evaluate the appropriateness of each model for the wheat yield data using residual diagnostics. Each model was fit using the full data set and estimated residuals were obtained in a manner similar to the process described in the previous paragraph for the GDEF model. Moran’s I is computed for each model to assess whether the estimated residuals had any spatial dependence and the Shapiro-Wilk test is applied to assess the normality of the residuals. The left side of table 4 contains the cross-validated average RMSEs, MAEs, coverage rates, and prediction interval widths for each model, while the right side contains the p-values associated with the Moran’s and Shapiro-Wilk tests.
RMSE MAE Coverage PI Width Moran’s I Shapiro-Wilk
GDEF 0.410 0.331 0.904 1.450 $p=0.460$ $p=0.262$
Matérn 0.436 0.355 0.936 1.689 $p=0.000$ $p=0.028$
H&H 0.684 0.552 0.898 2.209 $p=0.000$ $p=0.934$
Std. CAR 0.421 0.340 0.940 1.688 $p=0.000$ $p=0.002$
CAR 2-param. 0.422 0.342 0.936 1.647 $p=0.000$ $p=0.008$
Table 4: Comparison of predictive metrics and results from assumption tests across model. In addition to having the smallest out-of-sample prediction error, the GDEF model is the only one that produces estimated residuals that exhibit no spatial correlation.
The GDEF model has the lowest prediction error among models considered. In addition to this, it is the only model for which the Moran’s I statistic results in a failure to reject the null hypothesis of no spatial dependence in the model’s residuals. In other words, the GDEF model is the only model among these options with a covariance structure that is capable of fully describing the spatial dependence pattern present in the wheat yield data. Among models compared, the GDEF model produced the narrowest prediction intervals, though the coverage rate of these intervals appears to be somewhat under target, suggesting that the estimated marginal variance of predictions from the GDEF model may be too low; further investigation of this could be warranted. Despite considerably greater model complexity than the alternatives, the GDEF model’s out-of-sample predictive performance suggests that it is not overfit to these data and represents an improvement over simpler, more standard methodologies. While it could be argued that the differences in predictive performance between the GDEF and other models is not dramatic, it seems clear that the underlying covariance structure of the wheat yield data is most fully captured by our proposed methodology. We find this to be notable given the dataset’s history and reputation as a challenging spatial modeling problem.